# Bridge multisections of knotted surfaces in $S^{4}$

### v1

@@ 4.2. Band surgery
First, we make sure that no consecutive arcs are of type 1. Suppose, without loss of generality, that $\rho_{1}$, $\rho_{2}$, and $\rho_{3}$ are type 0, 1, and 1, respectively. Let $(\rho^{\prime},f^{\prime})$ be the tuple obtained by replacing $\rho_{2}$ from $\rho$ with a type 0 arc. We will see that $F_{\mathcal{T}[\rho,f]}$ can be obtained from $F_{\mathcal{T}[\rho^{\prime},f^{\prime}]}$ by adding a 1–handle or (at most) one unknotted $RP^{2}$ summand. The first case we study is when $\rho_{2}\cup\rho_{3}$ connects distinct components of $L_{2}$. Near the disks $\Delta_{1}$ and $\Delta_{2}$, $F_{\mathcal{T}[\rho,f]}$ looks like Part (d) of Figure 11. From Parts (a) and (b) of the same figure, we observe that $F_{\mathcal{T}[\rho,f]}$ is obtained by adding one 1–handle to $F_{\mathcal{T}[\rho^{\prime},f^{\prime}]}$. Suppose now that $\rho_{i}\cup\rho_{i+1}$ connects the same component of $L_{i}$ as in Figure 12(c). Here, parts (d) and (e) depict the surgered surface obtained by choosing the framing $f_{2}+f_{3}$ to be 0 and ±$\pm 1$, respectively. One can see that changing the framing $f_{2}+f_{3}$ from zero to ±$\pm 1$ amounts to replacing a subdisk of $\mathcal{D}_{1}\cup\mathcal{D}_{2}$ with a Möbius band. Thus, the surface $F_{\mathcal{T}[\rho,f]}$ with $f_{2}+f_{3}=\pm 1$ is obtained by connect summing an unknotted $RP^{2}$ to the surface represented by $\mathcal{T}[\rho,f]$ with $f_{2}+f_{3}=0$. In other words, we can assume that $f_{2}+f_{3}=0$. Now, from Parts (b) and (d) of Figure 12, we see that $F_{\mathcal{T}[\rho,f]}$ is isotopic to $F_{\mathcal{T}[\rho^{\prime},f^{\prime}]}$ via an isotopy supported in a neighborhood of $\Delta_{2}$.
@@
[…]
$\cdots 01\overset{\pm}{=}1X\cdots$$\overset{\#RP^{2}}{\longleftarrow}\cdots 001X\cdots$
[…]
In practice, if the first tangle in our tuple $(\rho,f)$ is type 0, we can apply these reductions to the corresponding string to get a world in $\{0,1\}$ with no consecutive ones. Here, by Example 4.7, this final vector corresponds to $(t-1)$ 1–handles being added to $F_{\mathcal{T}}$.
We run the counting process in Remark 4.10 with the admissible tuples in Figure 25. From $\mathcal{T}_{3}$ to $\mathcal{T}_{2}$ we gain one crosscap as the sequence of reductions looks as follows,
[…]
@@
[…]
@@ 5. Multiplane moves
The first three moves are the multiplane versions of the triplane moves from [MZ17]. Fix a $b$–bridge multiplane diagram $\mathcal{T}$. An interior Reidemeister move is an isotopy of the tangles of $\mathcal{T}$ that fixes the punctures. A mutual braid move is the result of appending the same $2b$–stranded braid to each tangle of $\mathcal{T}$. These correspond to ambient isotopies supported in a neighborhood of the central bridge surface. A multiple-sector perturbation is the result of dragging the multisected surface $F$ through certain disk $\Delta$ embedded in multiple consecutive 4-dimensional sectors. This move is a particular band surgery of $\mathcal{T}$ from Section 4.2 and will be discussed in detail in Section 5.1.
The last two types of multiplane moves, split and merge, change the number of sectors in a bridge multisection. They will be discussed in detail in Section 5.2. In Section 5.3, we discuss a particular combination of the perturbations and merge moves that allow us to obtain a triplane diagram from any given multiplane diagram. This combination is key in proving our uniqueness statement for bridge multisections (Thm 5.8).
@@ 5.1. Multiple-sector perturbations
Let $\mathcal{T}=(T_{1},\dots,T_{n})$ be a multiplane diagram for an embedded surface $F\subset S^{4}$. We say that a multiplane diagram $\mathcal{T}^{\prime}$ is a multiple-sector perturbation of $\mathcal{T}$ if $\mathcal{T}^{\prime}=\mathcal{T}[\rho,f]$ for an admissible tuple of framed arcs $(\rho,f)$ satisfying the following conditions for some integer $20\leq m\leq n-$ 2
@@
In this case, we may also refer to $\mathcal{T}^{\prime}$ as an $m$–sector perturbation of $\mathcal{T}$.
@@
In Lemma 5.1, we push our surface $F$ along a disk in the first $m$ sectors, this increases the bridge and patch numbers ($b$ and $c_{i}$’s) by one except for $c_{m+1}$ and $c_{n}$. If $n=3$ and $m=1$, i.e. a classic trisection perturbation, this process yields a $(b+1;c_{1},c_{2},c_{3}+1)$–bridge trisection [MZ17, Lem 6.1]. Figures 14 and 15 have examples of admissible arcs inducing two-sector perturbations.
@@
Notice that $m$–sector perturbations are not defined for $m=n-1,n$. One can interpret an $m$–perturbation as transferring saddles of $F$ from the sectors $\mathcal{D}_{1},\dots,\mathcal{D}_{m}$ to the rest of the sectors. Thus intuitively, if $m=n-1$, then the resulting surface $F^{\prime}\cap X_{n}$ may not be a trivial disk system. If $m=n$, ‘perturbing’ along a disk $\Delta\subset X_{1}\cup\dots\cup X_{n}$ can be interpreted as a compression of $F$ along $\Delta$, thus changing knotted surface.
@@
0–Sector perturbations change the multisected surface by an isotopy supported in a neighborhood of one 3-ball $B_{i}$. They are determined by one type 1 arc regardless of the framing. Following the definition of band surgery, one can see that the set of non-trivial links appearing as a subset of a cross-section of $\mathcal{T}$ is not affected by 0–sector perturbations. Although they are defined, 0–sector perturbations are redundant: if $\mathcal{T}^{\prime}$ is obtained from $\mathcal{T}$ by a 0-sector perturbation then $\mathcal{T}^{\prime}$ is also obtained from $\mathcal{T}$ by a $(n-2)$–sector perturbation.
@@ 5.2. Split and Merge
We now translate the merging criterion in Lemma 5.4 for merging a multisection in terms of dual bands. This may be used in practice for multiplane diagrams with many crossings. Recall that, by Lemma 2.4, consecutive tangles in a multiplane diagram are related by surgery along dual bands. In particular, if three consecutive tangles of $\mathcal{T}$ satisfy $T_{1}[v_{1}]=T_{2}$ and $T_{2}[v_{2}]=T_{3}$, then we can draw the bands $v_{2}$ in the same 3-ball as $T_{1}$ and $v_{1}$. If $v_{2}$ and $v_{1}$ share some endpoints, we choose to draw $v_{2}$ closer to the bridge surface than $v_{1}$ as in the top left of Figure 20. We drew concrete examples of such bands in Figure 19.
@@
The fact that $v=v_{1}\cup v_{2}$ is dual to $T_{1}$ implies that $T_{1}\cup\overline{T}_{3}$ is a $(b-|v|)$–component unlink and so the triplet $\mathcal{T}_{0}=(T_{1},T_{2},T_{3})$ is a $(b;b-|v_{1}|,b-|v_{2}|,b-|v|)$–triplane diagram. Moreover, after braid and interior Reidemeister moves, the tangle and the bands $T_{1}\cup v$ can be crossingless as in Figure 5. Thus, $T_{2}=T_{1}[v_{1}]$ and $T_{3}=T_{1}[v]$ are also crossingless tangles and so $\mathcal{T}_{0}$ is completely decomposable with underlying surface $F_{\mathcal{T}_{0}}$. As the Euler characteristic of $F_{\mathcal{T}_{0}}$ is $2(b-|v|)$, $F_{\mathcal{T}_{0}}$ must be a $(b-|v|)$–component unlink of unknotted 2–spheres. Hence, by Lemma 5.4, $\mathcal{T}^{\prime}$ is a multiplane diagram with $F_{\mathcal{T}^{\prime}}$ isotopic to $F_{\mathcal{T}}$. ∎
@@
$\mathcal{T}^{\prime}=\left(T^{(0)}_{1},\dots,T^{(m_{1}-1)}_{1},T^{(0)}_{2},\dots,T^{(m_{2}-1)}_{1},T^{(0)}_{3},\dots,T^{(m_{3}-1)}_{1}\right)$
@@ 5.3. Perturbation-merging combination
Figure 14. Turning a $(2;1)$–bridge 4–section into a $(3;1)$–trisection of an unknotted torus.
Figure 15. Turning a $(2;1)$–bridge 4–section into a $(3;1)$–trisection of a Klein bottle.
@@
Let $n\geq 4$ and let $\mathcal{T}=(T_{1},T_{2},T_{3},\dots,T_{n})$ be a $(b;c_{1},c_{2},\dots,c_{n})$–multiplane diagram. There exist $(b-c_{2})$ two-sector perturbations such that the resulting multiplane diagram $(T^{\prime}_{1},T^{\prime}_{2},T^{\prime}_{3},\dots,T^{\prime}_{n})$ can be merged into a multisection with $(n-1)$ sectors $(T^{\prime}_{1},T^{\prime}_{3},T^{\prime}_{4},\dots,T^{\prime}_{n})$ with invariants equal to
@@
The figures referenced in this proof depict the case $c_{2}=|T_{2}\cup\overline{T}_{3}|=1$; that being said, all of our arguments carry over to the case $c_{2}\geq 2$. After mutual braid moves and interior Reidemeister moves, we can represent $T_{2}$ and $T_{3}$ by a crossingless diagram as in Figure 16. In this figure, the dotted arcs (with blackboard framing) form admissible tuples for $\mathcal{T}$. Thus, they induce a sequence of $(b-c_{2})$ two-sector perturbations along disks in $X_{1}\cup X_{2}$. We denote by $\mathcal{T}^{\prime}=(T_{1}^{\prime},\dots,T_{n}^{\prime})$ the resulting bridge multisection. By Lemma 5.1, the invariants of $\mathcal{T}^{\prime}$ are given by
@@
We observe in Figure 17 that the link $T^{\prime}_{1}\cup\overline{T}^{\prime}_{3}$ is isotopic to $T_{1}\cup\overline{T}_{2}$. Thus, $\mathcal{T}_{0}=(T^{\prime}_{1},T^{\prime}_{2},T^{\prime}_{3})$ is a $(2b-c_{2};c_{1}+b-c_{2},b,c_{1})$triplane diagram. Via an Euler characteristic computation, we see that the underlying surface for $\mathcal{T}_{0}$ is a $c_{1}$–component link of 2–spheres. To apply Lemma 5.4, we must verify that $\mathcal{T}_{0}$ is completely decomposable. In Figure 18 we show a sequence of $(b-c_{2})$ 1-sector deperturbations that turn $\mathcal{T}_{0}$ into a $(b;c_{1},b,c_{1})$–bridge trisection. By Proposition 3.9, the resulting trisection is completely decomposable and so is $\mathcal{T}_{0}$. Lemma 5.4 implies then that $(T_{1},T_{3},T_{4},\dots,T_{n})$ is a multiplane diagram with underlying surface $F_{\mathcal{T}}$, as desired.
@@
We now explain the perturbation-merging process in Proposition 5.7 using bands instead of standard pairs. This may be useful in practice as dual bands can guide the necessary 2–sector perturbations. Figure 19 is an example of such a procedure. In there, we turn the $(3;2)$–bridge 4–sections of $m$–twist spun 2–bridge knots from Proposition 3.11 into the $(3;2)$–bridge trisections described by Meier and Zupan in [MZ17].
Figure 19. Turning a $(3;2)$–bridge 4–section into a $(4;2)$–trisection of an $m$–twist spun knot.
@@
Let $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ be two multiplane diagrams representing isotopic surfaces in 4–space. There is a finite sequence of multiplane moves that turns $\mathcal{T}_{1}$ into $\mathcal{T}_{2}$.

### v2

@@ 4.2. Band surgery
First, we make sure that no consecutive [^3] We consider the indices $n$ and $1$ to be consecutive as the tangles $T_{n}\cup\overline{T}_{n}$ bound disks in the multisected surface; see Definition 3.1. arcs are of type 1. Suppose, without loss of generality, that $\rho_{1}$, $\rho_{2}$, and $\rho_{3}$ are type 0, 1, and 1, respectively. Let $(\rho^{\prime},f^{\prime})$ be the tuple obtained by replacing $\rho_{2}$ from $\rho$ with a type 0 arc. We will see that $F_{\mathcal{T}[\rho,f]}$ can be obtained from $F_{\mathcal{T}[\rho^{\prime},f^{\prime}]}$ by adding a 1–handle or (at most) one unknotted $RP^{2}$ summand. The first case we study is when $\rho_{2}\cup\rho_{3}$ connects distinct components of $L_{2}$. Near the disks $\Delta_{1}$ and $\Delta_{2}$, $F_{\mathcal{T}[\rho,f]}$ looks like Part (d) of Figure 11. From Parts (a) and (b) of the same figure, we observe that $F_{\mathcal{T}[\rho,f]}$ is obtained by adding one 1–handle to $F_{\mathcal{T}[\rho^{\prime},f^{\prime}]}$. Suppose now that $\rho_{i}\cup\rho_{i+1}$ connects the same component of $L_{i}$ as in Figure 12(c). Here, parts (d) and (e) depict the surgered surface obtained by choosing the framing $f_{2}+f_{3}$ to be 0 and $\pm 1$, respectively. One can see that changing the framing $f_{2}+f_{3}$ from zero to $\pm 1$ amounts to replacing a subdisk of $\mathcal{D}_{1}\cup\mathcal{D}_{2}$ with a Möbius band. Thus, the surface $F_{\mathcal{T}[\rho,f]}$ with $f_{2}+f_{3}=\pm 1$ is obtained by connect summing an unknotted $RP^{2}$ to the surface represented by $\mathcal{T}[\rho,f]$ with $f_{2}+f_{3}=0$. In other words, we can assume that $f_{2}+f_{3}=0$. Now, from Parts (b) and (d) of Figure 12, we see that $F_{\mathcal{T}[\rho,f]}$ is isotopic to $F_{\mathcal{T}[\rho^{\prime},f^{\prime}]}$ via an isotopy supported in a neighborhood of $\Delta_{2}$.
@@
(1) $\cdots 10\cdots 01\cdots$ $\overset{\text{1--h}}{\longleftarrow}\cdots 10\cdots 00\cdots$
(2) $\cdots 01\neq 1\cdots$ $\overset{\text{1--h}}{\longleftarrow}\cdots 001\cdots$
(3) $\cdots 01\overset{\pm}{=}1\cdots$ $\overset{\#RP^{2}}{\longleftarrow}\cdots 001\cdots$
(4) $\cdots 01\overset{0}{=}1\cdots$ $\overset{\text{isotopy}}{\longleftarrow}\cdots 001\cdots$
In Rule (1), two non-consecutive type 1 arcs are separated by a string of some type zero arcs. The rule states that we can turn one of the type 1 arcs into a type 0 at the expense of removing a 1–handle from the multisected surface; this can be seen in Figure 13. Rule (2) turns the substring $01\neq 1$ into $001$ at the expense of removing a 1–handle; this can be seen in Figure 11. Rules (3) and (4) explain how to replace the substring $01\overset{\varepsilon}{=}1$ with $001$ at the expense of removing a crosscap summand or without altering the isotopy class of the multisected surface; this is observed in Figure 12 depending on whether $\varepsilon=\pm 1$ or 0.
In practice, if the first tangle in our tuple $(\rho,f)$ is type 0, we can apply Rules (1)–(4) to the corresponding string to get the word $100\cdots 0$. Here, one can see that for such a string, $\mathcal{T}[\rho,f]$ and $\mathcal{T}$ represent isotopic surfaces.
Example 4.12 (Tubings).
The admissible tuple in Example 4.7 is described by the string
$1\neq 1\neq\cdots\neq 10\cdots 0$
with $(m+1)$ consecutive 1’s and (nm1)(n-m-1) 0’s. Modulo cyclic permutation, this string can be reduced using Rule (2) to the string $100\cdots 0$. Thus, the band surgery described in Example 4.7 changes the surface by $m$ 1–handle additions.
Example 4.13 (More tubings).
The admissible tuple in Example 4.8 is described by the string
$0\cdots 010\cdots 010\cdots 010\cdots$
where any pair of 1’s is separated by at least one 0. Modulo cyclic permutation, this string can be reduced using Rule (1) to the string $100\cdots 0$. Thus, the band surgery described in Example 4.8 changes the surface by as many 1–handle additions as one less than the number of type one arcs in $\rho$.
Example 4.14 (Crosscap sums).
The admissible tuple in Example 4.9 is described by the string
$1\overset{\pm}{=}1\overset{\pm}{=}\cdots\overset{\pm}{=}10\cdots 0$
with $(m+1)$ consecutive 1’s and (nm1)(n-m-1) 0’s. Modulo cyclic permutation, this string can be reduced using Rule (3) to the string $100\cdots 0$. Thus, the band surgery described in Example 4.9 changes the surface by $m$ crosscap sums.
Example 4.15.
We discuss how the admissible tuples in Figure 25 change the multisected surfaces. From $\mathcal{T}_{3}$ to $\mathcal{T}_{2}$ we gain one crosscap as the sequence of reductions looks as follows,
$01\overset{0}{=}1\overset{\pm}{=}1\overset{0}{=}1\overset{\text{isotopy}}{\longleftarrow}001\overset{\pm}{=}1\overset{0}{=}1\overset{\#RP^{2}}{\longleftarrow}0001\overset{0}{=}1\overset{\text{isotopy}}{\longleftarrow}00001.$
@@
$01\neq 1\overset{\pm}{=}1\neq 1\overset{\text{1-h}}{\longleftarrow}001\overset{\pm}{=}1\neq 1\overset{\#RP^{2}}{\longleftarrow}0001\neq 1\overset{\text{1-h}}{\longleftarrow}00001.$
@@ 5. Multiplane moves
The first three moves are the multiplane versions of the triplane moves from [MZ17]. Fix a bb–bridge multiplane diagram $\mathcal{T}$. An interior Reidemeister move is an isotopy of the tangles of $\mathcal{T}$ that fixes the punctures. A mutual braid move is the result of appending the same 2b2b–stranded braid to each tangle of $\mathcal{T}$. These correspond to ambient isotopies supported in a neighborhood of the central bridge surface. A multiple-sector perturbation is the result of dragging the multisected surface $F$ through certain disk $\Delta$ embedded in multiple consecutive 4-dimensional sectors. This move is a particular band surgery of $\mathcal{T}$ from Section 4.2 and will be discussed in detail in Section 5.1.
The last two types of multiplane moves, split and merge, change the number of sectors in a bridge multisection. They will be discussed in detail in Section 5.2. In Section 5.3, we discuss a particular combination of the perturbations and merge moves that allow us to obtain a triplane diagram from any given multiplane diagram. This combination is key in proving our uniqueness statement for bridge multisections (Thm 5.10).
@@ 5.1. Multiple-sector perturbations
Let $\mathcal{T}=(T_{1},\dots,T_{n})$ be a multiplane diagram for an embedded surface $F\subset S^{4}$. We say that a multiplane diagram $\mathcal{T}^{\prime}$ is a multiple-sector perturbation of $\mathcal{T}$ if $\mathcal{T}^{\prime}=\mathcal{T}[\rho,f]$ for an admissible tuple of framed arcs $(\rho,f)$ satisfying the following conditions for some integer $0\leq m\leq n-2$
@@
We may also refer to $\mathcal{T}^{\prime}$ as an mm–sector perturbation of $\mathcal{T}$. Observe that mm–sector perturbation corresponds to the band surgery in Example 4.7 with $(m,r)=(m,0)$.
@@
In Lemma 5.1, we push our surface $F$ along a disk in the first $m$ sectors, increasing the bridge and patch numbers ($b$ and $c_{i}$’s) by one, except for $c_{m+1}$ and $c_{n}$. If $n=3$ and $m=1$, i.e. a classic trisection perturbation, this process yields a $(b+1;c_{1}+1,c_{2},c_{3})$–bridge trisection [MZ17, Lem 6.1]. Figures 14 and 15 have examples of admissible arcs inducing two-sector perturbations.
@@
Notice that mm–sector perturbations are not defined for m=n1,nm=n-1,n. One can interpret an mm–perturbation as transferring saddles of $F$ from the sectors $\mathcal{D}_{1},\dots,\mathcal{D}_{m}$ to the rest of the sectors. Intuitively, if m=n1m=n-1, then the resulting surface $F^{\prime}\cap X_{n}$ may not be a trivial disk system. If $m=n$, ‘perturbing’ along a disk $\Delta\subset X_{1}\cup\dots\cup X_{n}$ can be interpreted as a compression of $F$ along $\Delta$, thus changing the knotted surface.
@@
0–Sector perturbations change the multisected surface by an isotopy supported in a neighborhood of one 3–ball $B_{i}$. They are determined by one type 1 arc regardless of the framing. Following the definition of band surgery, one can see that the set of non-trivial links appearing as a subset of a cross-section of $\mathcal{T}$ is not affected by 0–sector perturbations. Although they are defined, 0–sector perturbations are redundant: if $\mathcal{T}^{\prime}$ is obtained from $\mathcal{T}$ by a 0-sector perturbation then $\mathcal{T}^{\prime}$ is also obtained from $\mathcal{T}$ by a (n2)(n-2)–sector perturbation.
@@ 5.2. Split and Merge
We now translate the merging criterion in Lemma 5.4 for merging a multisection in terms of dual bands. This may be used in practice for multiplane diagrams with many crossings. Recall that, by Lemma 2.5, consecutive tangles in a multiplane diagram are related by surgery along dual bands. In particular, if three consecutive tangles of $\mathcal{T}$ satisfy $T_{1}[v_{1}]=T_{2}$ and $T_{2}[v_{2}]=T_{3}$, then we can draw the bands $v_{2}$ in the same 3–ball as $T_{1}$ and $v_{1}$. If $v_{2}$ and $v_{1}$ share some endpoints, we choose to draw $v_{2}$ closer to the bridge surface than $v_{1}$ as in the top left of Figure 20. We drew concrete examples of such bands in Figure 19.
@@
The fact that $v=v_{1}\cup v_{2}$ is dual to $T_{1}$ implies that $T_{1}\cup\overline{T}_{3}$ is a (b|v|)(b-|v|)–component unlink and so the triplet $\mathcal{T}_{0}=(T_{1},T_{2},T_{3})$ is a $(b;b-|v_{1}|,b-|v_{2}|,b-|v|)$–triplane diagram. Moreover, after braid and interior Reidemeister moves, the tangle and the bands $T_{1}\cup v$ can be crossingless as in Figure 5. Thus, $T_{2}=T_{1}[v_{1}]$ and $T_{3}=T_{1}[v]$ are also crossingless tangles and so $\mathcal{T}_{0}$ is completely decomposable with underlying surface $F_{\mathcal{T}_{0}}$. As the Euler characteristic of $F_{\mathcal{T}_{0}}$ is 2(b|v|)2(b-|v|), $F_{\mathcal{T}_{0}}$ must be a (b|v|)(b-|v|)–component unlink of unknotted 2–spheres. Hence, by Lemma 5.4, $\mathcal{T}^{\prime}$ is a multiplane diagram with $F_{\mathcal{T}^{\prime}}$ isotopic to $F_{\mathcal{T}}$. ∎
@@
$\mathcal{T}^{\prime}=\left(T^{(0)}_{1},\dots,T^{(m_{1}-1)}_{1},T^{(0)}_{2},\dots,T^{(m_{2}-1)}_{2},T^{(0)}_{3},\dots,T^{(m_{3}-1)}_{3}\right)$
@@
Remark 5.7 (Motivation).
Thin multisections permit the study of knotted surfaces as loops in certain complexes. Some of Engelhardt’s PhD dissertation discusses how loops in the pants complex of planar surfaces codify the same information as a thin bridge multisection. For closed 4-manifolds, the study of thin multisections was carried out by Islambouli in [ISL21].
In light of Theorem 5.6, one can study bounds on the number of sectors needed to achieve a thin bridge multisection of a given knotted surface. If $n=3$, it follows from Proposition 3.11 that thin bridge trisections represent unlinks of unknotted 2-spheres and at most one unknotted projective plane. The unknotted torus admits a thin bridge 4–section; see Figure 14.
Problem 5.8.
Compute the smallest $n\geq 3$ such that $F\subset S^{4}$ admits a thin bridge nn–section.
@@ 5.3. Perturbation-merging combination
Figure 14. Turning a (2;1)(2;1)–bridge 4–section into a (3;1)(3;1)–trisection of an unknotted torus.
Figure 15. Turning a (2;1)(2;1)–bridge 4–section into a (3;1)(3;1)–trisection of a Klein bottle.
@@
Let $n\geq 4$ and let $\mathcal{T}=(T_{1},T_{2},T_{3},\dots,T_{n})$ be a $(b;c_{1},c_{2},\dots,c_{n})$–multiplane diagram. There exist $(b-c_{2})$ two-sector perturbations such that the resulting multiplane diagram $(T^{\prime}_{1},T^{\prime}_{2},T^{\prime}_{3},\dots,T^{\prime}_{n})$ can be merged into a multisection with (n1)(n-1) sectors $(T^{\prime}_{1},T^{\prime}_{3},T^{\prime}_{4},\dots,T^{\prime}_{n})$ with invariants equal to
@@
The figures referenced in this proof depict the case $c_{2}=|T_{2}\cup\overline{T}_{3}|=1$; that being said, all of our arguments carry over to the case $c_{2}\geq 2$. After mutual braid moves and interior Reidemeister moves, we can represent $T_{2}$ and $T_{3}$ by a crossingless diagram as in Figure 16. In this figure, the arcs (with blackboard framing) form admissible tuples for $\mathcal{T}$. Thus, they induce a sequence of $(b-c_{2})$ two-sector perturbations along disks in $X_{1}\cup X_{2}$. We denote by $\mathcal{T}^{\prime}=(T_{1}^{\prime},\dots,T_{n}^{\prime})$ the resulting bridge multisection. By Lemma 5.1, the invariants of $\mathcal{T}^{\prime}$ are given by the new tuple
@@
We observe in Figure 17 that the link $T^{\prime}_{1}\cup\overline{T}^{\prime}_{3}$ is isotopic to $T_{1}\cup\overline{T}_{2}$. Thus, $\mathcal{T}_{0}=(T^{\prime}_{1},T^{\prime}_{2},T^{\prime}_{3})$ is a $(2b-c_{2};c_{1}+b-c_{2},b,c_{1})$triplane diagram. Via an Euler characteristic computation, we see that the underlying surface for $\mathcal{T}_{0}$ is a $c_{1}$–component link of 2–spheres. To apply Lemma 5.4, we must verify that $\mathcal{T}_{0}$ is completely decomposable. In Figure 18 we show a sequence of $(b-c_{2})$ 1-sector deperturbations that turn $\mathcal{T}_{0}$ into a $(b;c_{1},b,c_{1})$–bridge trisection. By Proposition 3.10, the resulting trisection is completely decomposable and so is $\mathcal{T}_{0}$. Lemma 5.4 implies then that $(T^{\prime}_{1},T^{\prime}_{3},T^{\prime}_{4},\dots,T^{\prime}_{n})$ is a multiplane diagram with underlying surface $F_{\mathcal{T}}$, as desired.
@@
We now explain the perturbation-merging process in Proposition 5.9 using bands instead of standard pairs. This may be useful in practice as dual bands can guide the necessary 2–sector perturbations. Figure 19 is an example of such a procedure. In there, we turn the (3;2)(3;2)–bridge 4–sections of mm–twist spun 2–bridge knots from Proposition 3.12 into the (4;2)(4;2)–bridge trisections described by Meier and Zupan in [MZ17].
Figure 19. Turning a (3;2)(3;2)–bridge 4–section into a (4;2)(4;2)–trisection of an mm–twist spun knot.
@@
Let $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ be two multiplane diagrams representing isotopic surfaces in $S^{4}$. There is a finite sequence of multiplane moves that turns $\mathcal{T}_{1}$ into $\mathcal{T}_{2}$.